OpenStudy (anonymous):

a theater has 43 seats in the first row, 45 seats in the second row, 47 seats in the third row, and so on. is it arithmetic or geometric? a^n=? there are ? seats in row 17?

OpenStudy (anonymous):

it's arithmetic because the terms have a common difference (2, in this case).

OpenStudy (anonymous):

i never saw a question like this before i do not understand

OpenStudy (anonymous):

so how do i solve for a^n= and how manyseats

OpenStudy (anonymous):

i ask my substitute teacher and she dont know ??? clueless

OpenStudy (anonymous):

Does the problem tell you how many total rows there are?

OpenStudy (anonymous):

it says how many seats are there in row 17?

OpenStudy (anonymous):

write a general term for sequence a^n

OpenStudy (anonymous):

Row 3 has 47 seats, so Row 17 has 47 + (14*2). that is the third row, plus fourteen times (17-3) the difference between the rows. So Row 17 has 75 seats.

OpenStudy (anonymous):

The formula for the general term is: a1 + (n-1) d this is the first term + (the term number - 1)*difference

OpenStudy (anonymous):

ok so for a^n= 47+(14*2) for the term

OpenStudy (anonymous):

For row 17, = 43 +(17-1)*2

OpenStudy (anonymous):

which is just plugging the numbers into the an=a1+(n-1)d also it's not a^n, remember the n is subscript, not like a power. :)

OpenStudy (anonymous):

ok i see how you gather the information to get 17 =75 seats

OpenStudy (anonymous):

thank you and i am still veiwing this formula

OpenStudy (anonymous):

No problem. Keep the formula in mind as it's going to come useful all the time with sequences.

OpenStudy (anonymous):

i need this as a note to show to the class becuase this substitute teacher slow